OpenStudy (anonymous):

Simplify the following expression. 2x − 2y + 3z − 2x − y + 2z = ? 4x − 3y + 5z −2y − 3y + 5z −2x − 3y + 6z −3y + 5z

OpenStudy (anonymous):

combine like terms. 2x-2x = 0x -2y -y = -3y 3z + 2z = 5z. -3y+5z